OpenStudy (anonymous):

Roger swam 400 meters in 3 minutes and 47.2 seconds. was his average speed for his swim more than 2 meters per second?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Roger is a tank.

OpenStudy (anonymous):

3 minutes and 47.2 seconds = (3*60)+47.2 seconds = 180+47.2 = 227.2 400 meters / 227.2 seconds = 1.76 meters per second

OpenStudy (anonymous):

1.76 m/s

OpenStudy (anonymous):

formula for average speed is S=D/T